# Foundation Models in Smart Agriculture: Basics, Opportunities, and Challenges

### 21 Dec 2023 v3

@@ 3.2.2 VFMs in agriculture
Vision Foundation Models (VFMs, as detailed in Section 2.2.2), encompassing algorithms for core computer vision tasks like classification, detection, segmentation, and generative modeling, have revolutionized traditional computer vision methodologies. Trained on extensive image datasets, these VFMs can be efficiently adapted to novel domains, such as agriculture, with minimal fine-tuning and a limited set of labeled images. Recently, the agricultural community has shown a growing interest in VFMs, as seen in recent studies (Yang et al., 2023c, Williams et al., 2023).
In Yang et al. (2023c), the Segment Anything Model (SAM, Section  2.2.2) was employed for chicken segmentation tasks in a zero-shot manner using both part-based segmentation and the use of infrared thermal images (Figure 3). Experimental results showed that SAM outperformed the other VFMs algorithms, such as SegFormer and SETR, in both whole and partial chicken segmentation. For example, SAM achieved a mean Intersection over Union (mIoU) of 94.80% in the chicken segmentation tasks, while SegFormer and SETR only achieved mIoUs of 43.22% and 42.90%, which greatly shows the effectiveness of SAM on the poultry segmentation tasks. In addition, a novel single-bird tracking algorithm, integrating SAM, YOLOX (Ge et al., 2021) and ByteTracker (Zhang et al., 2022b), was proposed for tracking the activities for individual birds across video frames by extracting the location information (i.e., bounding boxes provided by SAM) in each image and then tracking the birds over time with YOLOX and ByteTracker. Despite its promising performance, the tracking algorithm lacked end-to-end capability. The object tracking FMs (Yang et al., 2023a) showcases impressive performance in video object tracking and segmentation, requiring very little human participation, thereby holding the potential to serve as an end-to-end method.
In Williams et al. (2023), an automatic leaf segmentation pipeline, named “Leaf Only SAM”, was introduced for zero-shot segmentation of potato leaves. When compared with a fine-tuned Mask R-CNN (He et al., 2017) model on the annotated potato leaf dataset, Leaf Only SAM demonstrated an average recall and precision of 63.2% and 60.3%, respectively. These results are evidently lower than the Mask R-CNN model’s average recall and precision of 78.7% and 74.7%, respectively. However, it’s important to note that Leaf Only SAM does not necessitate any fine-tuning or an annotated dataset, thereby showing notable zero-shot generalization capabilities. These findings suggest that the direct application of FMs in a zero-shot manner may not always result in high performance due to potential distribution shifts (Chawla et al., 2021). This will be further discussed in Section  4.3.
@@ 3.2.3 MFMs in agriculture
Multimodal foundation models (MFMs, Section  2.2.3) introduce a novel class of artificial intelligence models. These leverage and incorporate various types of data inputs including text, video, audio, and images. The agricultural community has begun to delve into multimodal learning within agricultural applications (Bender et al., 2020, Garg et al., 2021, Parr et al., 2021, Cao et al., 2023a, Dhakshayani and Surendiran, 2023) to enhance performance through the utilization of multimodal data. For instance, Bender et al. (2020) open-sourced a multimodal dataset designed for agricultural robotics, which was gathered from cauliflower and broccoli fields, to facilitate robotics and machine learning research activity in agriculture. The dataset includes a variety of data such as stereo color, thermal and hyperspectral imagery, and crucial environmental details like weather and soil conditions. In Cao et al. (2023a), a multi-modal language model based on image-text-label information was proposed for cucumber disease recognition. They effectively merged label information across multiple domains by employing image-text multimodal contrastive learning and image self-supervised contrastive learning. This methodology helped measure the distance of samples within the common image-text-label space. The experimental results revealed that this novel approach attained a recognition accuracy rate of 94.84% on a modestly sized multimodal cucumber disease dataset.
MFMs have also been discussed in the agriculture domain. Lu et al. (2023) investigated the potential of GPT-4, a multimodal language and vision model (as referenced in ChatGPT, Section 2.2.1), in achieving Artificial General Intelligence (AGI). In Stella et al. (2023), Large Language Models (LLMs, e.g., ChatGPT) were utilized in the design phase of robotic systems for tomato picking. However, it’s noteworthy that current applications are predominantly based on text-image data and their use is often confined to question-answering tasks. Indeed, there is a conspicuous gap in applications related to agricultural robotics where inputs such as images, text, or voice (e.g., human instructions), and depth information (e.g., from LiDAR or laser sensors) are used. These robots, which are typically designed for tasks like fruit harvesting or crop monitoring (Tao and Zhou, 2017), could greatly benefit from the application of MFMs, which will be further discussed in Section  3.3.5.
@@ 3.2.4 RLFMs in agriculture
Deep reinforcement learning (DRL) (Arulkumaran et al., 2017, Henderson et al., 2018) is an emerging field that fuses deep learning with reinforcement learning to tackle decision-making tasks. Its impressive performance has been demonstrated across a broad spectrum of applications, from computer games (Mnih et al., 2013, 2015), to intelligent transportation systems (Chu et al., 2019, Chen et al., 2023), and robot control (Gu et al., 2017, Lee et al., 2020). Recently, the agricultural community has shown heightened interest in DRL (Binas et al., 2019, Gandhi, 2022), leading to a host of innovative applications such as crop management (Gautron et al., 2022, Din et al., 2022), agriculture robots (Yang et al., 2022b), and smart irrigation systems (Zhou, 2020). However, due to their reliance on learning tasks from scratch without comprehensive knowledge of vision, language, or other datasets, these methods often face challenges in terms of generalization and sample efficiency (Yang et al., 2023b).
Although there currently exist no published papers specifically focusing on the application of reinforcement learning foundation models (RLFMs, Section 2.2.4) in agricultural scenarios, the field nonetheless presents vast untapped opportunities for exploration and innovation. For instance, Gato (Reed et al., 2022) and AdA (Team et al., 2023), discussed in Section 2.2.4, exhibit promising potential for handling decision tasks that involve multimodal, multi-task contexts, paving the way for innovative agriculture applications in the future.
@@ 3.3 Emerging opportunities
In this subsection, we discuss some emerging areas in the agricultural space where these foundation models are most applicable.
@@ 3.3.1 Smart crop management
To achieve sustainable farming, smart crop management (Eli-Chukwu, 2019, Jaramillo et al., 2020) has evolved as a novel research direction. It integrates advanced methodologies from artificial intelligence, IoT, and big data into crop management tasks, including plant growth monitoring (Lakshmi and Gayathri, 2017), disease detection (Xu et al., 2022a), yield monitoring (Khanal et al., 2018, Kayad et al., 2019), and harvesting (Meshram et al., 2021). Yet, developing such algorithms for precision crop management is complex and challenging, particularly when considering factors like diverse crop varieties, environmental dynamics, and limited or hard-to-acquire training datasets. Furthermore, these systems often necessitate diverse data inputs sourced from varying domains. For example, while a deep learning method is developed on satellite imagery to assess crop health, it simultaneously might need soil sensor data to understand underground moisture levels and market data to predict the optimal time for harvest based on pricing trends. Designing such algorithms requires deep research insights, making it challenging to smoothly incorporate knowledge from unrelated fields.
In light of the aforementioned challenges, Foundation Models (FMs) offer a promising solution. These models, trained on datasets spanning multiple domains, can be seamlessly adapted to new areas with minimal fine-tuning. They hold significant potential as tools for smart crop management. For instance, Language Foundation Models (LFMs, Section  2.2.1) empower researchers to tap into knowledge beyond their personal expertise, facilitating software development and design (Stella et al., 2023). Moreover, multi-modal foundation models (MFMs, Section  2.2.3) integrate visual, textual, and sometimes auditory data, providing a comprehensive analytical lens, proving invaluable in deciphering intricate agricultural situations, notably in precision crop management.
@@ 3.3.2 Smart plant breeding
Smart plant breeding, a breakthrough approach, combines data from genotype, environment, and their interaction to optimize crop varieties (Xu et al., 2022b, Zhang et al., 2023c). Employing sophisticated “omics” technologies, artificial intelligence, and big data, it constructs effective models to predict plant traits and performance (phenotype). The objective of this “smart breeding” is to promote genetic gains, shorten breeding cycles, and develop plants that are acclimatized to specific and evolving environments (Xu et al., 2022b). However, the accuracy in phenotype prediction is intrinsically tied to the processing of multidimensional or multimodal data, sourced from spatiotemporal omics, encompassing genotype, environment, and their interaction. Such a data-rich environment poses significant challenges to conventional machine learning methodologies (Xu et al., 2022b).
In response to this complexity, multimodal foundation models (MFMs, Section  2.2.3) may emerge as a potential solution. These models, designed to handle and learn from varied data types across multiple domains, could significantly improve the accuracy and efficiency of predicting plant phenotypes, thereby pushing the field of smart plant breeding toward unprecedented possibilities.
@@ 3.3.4 Precision aquaculture farming
Precision aquaculture farming (O’Donncha and Grant, 2019, Føre et al., 2018, Antonucci and Costa, 2020), also known as smart aquaculture or precision fish farming, is a modern approach to fish and shellfish farming that integrates advanced technologies like IoT and data-driven techniques to optimize production efficiency, environmental sustainability, and animal welfare. It is often characterized by hundreds of interconnected sensors that store and serve data to gather essential information such as temperature, oxygen, weather, chlorophyll, fish tags, imagery data, and 3D sensor data, which can be utilized for monitoring fish variables (species, health, etc.), behavioral analysis, feeding decisions, and water quality prediction (Yang et al., 2021). Recently, deep learning (DL) has emerged as a promising solution to address the problems in precision aquaculture farming. For example, in Måløy et al. (2019), a dual-stream recurrent network was adopted to classify fish feeding status for salmon, yielding 80% feeding classification accuracy. However, these DL approaches rely heavily on large-scale and labeled datasets for training (Yang et al., 2021). Underwater conditions can be adverse, with poor illumination and low visibility in turbid water, as well as cluttered backgrounds, making it difficult to acquire high-quality data. Besides, it is very challenging to integrate, handle, and process numerous and varied data across multiple domains.
@@
Fortunately, FMs have become an encouraging solution to facilitate the realization of precision aquaculture farming, representing the unique capability to handle diverse data from multiple domains through few-short or zero-short learning. For example, generation-based FMs can be harnessed for data creation, while RLFMs can be utilized for decision-making in fish farming. We firmly believe that the application of FMs could significantly enhance the accuracy and efficiency of associated applications in precision aquaculture farming.
3.3.5 Agriculture robots
@@
Agricultural robots, often referred to as “AgriRobots”, are at the forefront of transforming modern agriculture (Bechar and Vigneault, 2016). AgriRobots can automate various farming tasks such as planting, weeding, harvesting, and crop health monitoring (Fountas et al., 2020). For instance, AI-guided robots can identify and remove weeds with precision, significantly reducing the use for herbicides. Similarly, drones equipped with advanced imaging sensors can monitor fields for signs of disease or pests, enabling early intervention and reducing yield losses. Similar to the contexts of smart breeding and smart livestock farming, the development and implementation of AgriRobots require multidisciplinary knowledge including robotics, AI, plant science, environmental science, and data analytics. This mixed approach involves handling and processing a significant amount of data from various domains, making it a complex endeavor (Bechar and Vigneault, 2016, Fountas et al., 2020).
@@
However, these challenges can be guided effectively by the potential application of FMs. These models, trained on large-scale multimodal data, can provide the robust predictive capabilities, which are necessary for advanced farming operations. For instance, Reinforcement Learning Foundation Models (RLFM, Section  2.2.4) could enable precision control of AgriRobots (Reed et al., 2022, Team et al., 2023), while Language Foundation Models (LFM, Section  2.2.1) could provide valuable insights during the design process of AgriRobots (Stella et al., 2023). By incorporating insights from diverse domains, FMs can offer an integrated solution, driving the development of more effective and efficient agricultural robots, and thus, revolutionizing the future of agriculture.
@@ 4.1 Data collection
The process of data collection for training FMs in agricultural applications presents unique challenges due to the inherent variability and diversity in agriculture, influenced by factors such as crop types, growth stages, soil conditions, weather patterns, farming practices, etc. These factors result in a complex and highly diverse dataset, making its collection and standardization a challenging task (Lu and Young, 2020). Firstly, it is critical to ensure the collected data is accurate, comprehensive, and adequately captures the wide variability found in different farming environments or diverse range of growth phases. However, acquiring high-quality data can be time-consuming, labor-intensive, and expensive, especially considering the need for ground truth labeling in supervised learning scenarios (Li et al., 2023a) and the variety growth stages of plants or livestocks. To tackle these challenges, generative models (Lu et al., 2022) that create high-fidelity images and label-efficient learning methods to save on tedious labeling costs (Li et al., 2023a, Zhang et al., 2023e) show promises. Secondly, data privacy and ownership pose challenges as farms are typically privately owned. Farmers may be reluctant to share data due to concerns over privacy or potential commercial exploitation. Novel frameworks, such as federated learning (Zhang et al., 2021a), can address these concerns, paving the way for broader access to high-quality, diverse agricultural datasets. Temporal dynamics represent another significant hurdle. Agricultural processes inherently change over time, influenced by daily fluctuations, seasonal variations, or annual changes. This requires the collection of time-series data (e.g., data across years), introducing another layer of complexity. To handle temporal dynamics, techniques such as time-series analysis (Ostrom, 1990), dynamic models based on Long Short-Term Memory (LSTM) (Graves and Graves, 2012), and data augmentation (Lu et al., 2022) can be employed to better account for the inherent temporal dynamics in agricultural data.

### 17 Mar 2024 v4

@@ 3.2.2 VFMs in agriculture
VFMs (as detailed in Section  2.2.2), encompassing algorithms for core computer vision tasks like classification, detection, segmentation, and generative modeling, have revolutionized traditional computer vision methodologies. Trained on extensive image datasets, these VFMs can be efficiently adapted to novel domains, such as agriculture, with minimal fine-tuning and a limited set of labeled images. Recently, the agricultural community has shown a growing interest in VFMs, as seen in recent studies (Yang et al., 2023c, Williams et al., 2023).
In Yang et al. (2023c), SAM (Section  2.2.2) is employed for chicken segmentation tasks in a zero-shot manner using both part-based segmentation and the use of infrared thermal images (Figure 3 ). Experimental results show that SAM outperformed the other VFMs algorithms, such as SegFormer and SETR, in both whole and partial chicken segmentation. For example, SAM achieves a mean Intersection over Union (mIoU) of 94.80% in the chicken segmentation tasks, while SegFormer and SETR only achieve mIoUs of 43.22% and 42.90%, which greatly shows the effectiveness of SAM on the poultry segmentation tasks. In addition, a novel single-bird tracking algorithm, integrating SAM, YOLOX (Ge et al., 2021) and ByteTracker (Zhang et al., 2022b), is proposed for tracking the activities for individual birds across video frames by extracting the location information (i.e., bounding boxes provided by SAM) in each image and then tracking the birds over time with YOLOX and ByteTracker. Despite its promising performance, the tracking algorithm lacked end-to-end capability. The object tracking FMs (Yang et al., 2023a) show impressive performance in video object tracking and segmentation, requiring very little human participation, thereby holding the potential to serve as an end-to-end method.
In Williams et al. (2023), an automatic leaf segmentation pipeline, named “Leaf Only SAM”, is introduced for zero-shot segmentation of potato leaves. When compared with a fine-tuned Mask R-CNN (He et al., 2017) model on the annotated potato leaf dataset, Leaf Only SAM demonstrates an average recall and precision of 63.2% and 60.3%, respectively. These results are lower than the Mask R-CNN model’s average recall and precision of 78.7% and 74.7%, respectively. However, it’s important to note that Leaf Only SAM does not necessitate any fine-tuning or an annotated dataset, thereby showing notable zero-shot generalization capabilities. These findings suggest that the direct application of FMs in a zero-shot manner may not always result in high performance due to potential distribution shifts (Chawla et al., 2021). This will be further discussed in Section  4.3.
@@ 3.2.3 MFMs in agriculture
MFMs (Section  2.2.3) introduce a novel class of AI models. These leverage and incorporate various types of data inputs including text, video, audio, and images. The agricultural community has begun to delve into multimodal learning within agricultural applications (Bender et al., 2020, Garg et al., 2021, Parr et al., 2021, Cao et al., 2023a, Dhakshayani and Surendiran, 2023) to enhance performance through the utilization of multimodal data. For instance, Bender et al. (2020) open-sources a multimodal dataset designed for agricultural robotics, which is gathered from cauliflower and broccoli fields, to facilitate robotics and machine learning research activity in agriculture. The dataset includes a variety of data such as stereo color, thermal and hyperspectral imagery, and crucial environmental details like weather and soil conditions. In Cao et al. (2023a), a multi-modal language model based on image-text-label information is proposed for cucumber disease recognition. They effectively merge label information across multiple domains by employing image-text multimodal contrastive learning and image self-supervised contrastive learning. This methodology helps measure the distance of samples within the common image-text-label space. The experimental results reveal that this novel approach attains a recognition accuracy rate of 94.84% on a modestly sized multimodal cucumber disease dataset.
MFMs have also been discussed in the agriculture domain. Lu et al. (2023) investigates the potential of GPT-4, a multimodal language and vision model (as referenced in ChatGPT, Section 2.2.1), in achieving AGI. In Stella et al. (2023), LLMs (e.g., ChatGPT) are utilized in the design phase of robotic systems for tomato picking. In Tan et al. (2023), GPT-4 is applied for geographical, environmental, agricultural, and urban planning applications, such as crop type identification.
However, it is noteworthy that current applications are predominantly based on text-image data, and their use is often confined to question-answering tasks. Indeed, there is a conspicuous gap in applications related to agricultural robotics where inputs such as images, text, or voice (e.g., human instructions), and depth information (e.g., from LiDAR or laser sensors) are used. These robots, which are typically designed for tasks like fruit harvesting or crop monitoring (Tao and Zhou, 2017), could greatly benefit from the application of MFMs, which will be further discussed in Section  3.3.5.
@@ 3.2.4 RLFMs in agriculture
DRL (Arulkumaran et al., 2017, Henderson et al., 2018) is an emerging field that fuses deep learning with reinforcement learning to tackle decision-making tasks. Its impressive performance has been demonstrated across a broad spectrum of applications, from computer games (Mnih et al., 2013, 2015), to intelligent transportation systems (Chu et al., 2019, Chen et al., 2023), and robot control (Gu et al., 2017, Lee et al., 2020). Recently, the agricultural community has shown heightened interest in DRL (Binas et al., 2019, Gandhi, 2022), leading to a host of innovative applications such as crop management (Gautron et al., 2022, Din et al., 2022), agriculture robots (Yang et al., 2022b), and smart irrigation systems (Zhou, 2020). However, due to their reliance on learning tasks from scratch without comprehensive knowledge of vision, language, or other datasets, these methods often face challenges in terms of generalization and sample efficiency (Yang et al., 2023b).
Although there currently exist no published papers specifically focusing on the application of RLFMs (Section 2.2.4) in agricultural scenarios, the field nonetheless presents vast untapped opportunities for exploration and innovation. For instance, Gato (Reed et al., 2022) and AdA (Team et al., 2023a), discussed in Section  2.2.4, exhibit promising potential for handling decision tasks that involve multimodal, multi-task contexts, paving the way for innovative agriculture applications in the future.
Table 2 summarizes the applications of LLMs and FMs in agriculture.
Table 2: Applications of LLMs and FMs in agriculture.
Model Type Reference Problem Method Applications
LFMs Stella et al. (2023) Robotic picking ChatGPT Agriculture robots
Tzachor et al. (2023) Agricultural recommendations GPT Agricultural extension
Lu et al. (2023) Data processing GPT Organizing unstructured metadata
Yang et al. (2024) Plant science LLAMA Plant science-related tasks
Shutske (2023) Agricultural safety & health ChatGPT Agricultural safety and health questions
Kuska et al. Ai-Chatbots ChatGPT Plant disease management
VFMs Yang et al. (2023c) Segmentation and tracking SAM Chicken segmentation, bird tracking
Williams et al. (2023) Segmentation SAM Automatic leaf segmentation
MFMs Cao et al. (2023a) Disease recognition Image-text data Cucumber disease recognition
Stella et al. (2023) Robotic picking Image-text data Tomato picking
Tan et al. (2023) Remote Sensing GPT4 Crop type identification
RLFMs - - - -
@@ 3.3 Emerging opportunities
In this subsection, some emerging areas in the agricultural space are discussed where these FMs are most applicable.
@@ 3.3.1 Smart crop management
To achieve sustainable farming, smart crop management (Eli-Chukwu, 2019, Jaramillo et al., 2020) has evolved as a novel research direction. It integrates advanced methodologies from AI, IoT, and big data into crop management tasks, including plant growth monitoring (Lakshmi and Gayathri, 2017), disease detection (Xu et al., 2022b), yield monitoring (Khanal et al., 2018, Kayad et al., 2019), and harvesting (Meshram et al., 2021). Yet, developing such algorithms for precision crop management is complex and challenging, particularly when considering factors like diverse crop varieties, environmental dynamics, and limited or hard-to-acquire training datasets. Furthermore, these systems often necessitate diverse data inputs sourced from varying domains. For example, while a deep learning method is developed on satellite imagery to assess crop health, it simultaneously might need soil sensor data to understand underground moisture levels and market data to predict the optimal time for harvest based on pricing trends. Designing such algorithms requires deep research insights, making it challenging to smoothly incorporate knowledge from unrelated fields.
In light of the aforementioned challenges, FMs offer a promising solution. These models, trained on datasets spanning multiple domains, can be seamlessly adapted to new areas with minimal fine-tuning. They hold significant potential as tools for smart crop management. For instance, LFMs (Section  2.2.1) empower researchers to tap into knowledge beyond their personal expertise, facilitating software development and design (Stella et al., 2023). Moreover, MFMs (Section  2.2.3) integrate visual, textual, and sometimes auditory data, providing a comprehensive analytical lens, proving invaluable in deciphering intricate agricultural situations, notably in precision crop management.
@@ 3.3.2 Smart plant breeding
Smart plant breeding, a breakthrough approach, combines data from genotype, environment, and their interaction to optimize crop varieties (Xu et al., 2022c, Zhang et al., 2023d). Employing sophisticated “omics” technologies, AI, and big data, it constructs effective models to predict plant traits and performance (phenotype). The objective of this “smart breeding” is to promote genetic gains, shorten breeding cycles, and develop plants that are acclimatized to specific and evolving environments (Xu et al., 2022c). However, the accuracy in phenotype prediction is intrinsically tied to the processing of multidimensional or multimodal data, sourced from spatiotemporal omics, encompassing genotype, environment, and their interaction. Such a data-rich environment poses significant challenges to conventional machine learning methodologies (Xu et al., 2022c).
In response to this complexity, MFMs (Section  2.2.3) may emerge as a potential solution. These models, designed to handle and learn from varied data types across multiple domains, could significantly improve the accuracy and efficiency of predicting plant phenotypes, thereby pushing the field of smart plant breeding toward unprecedented possibilities.
@@ 3.3.4 Precision aquaculture farming
Precision aquaculture farming (O’Donncha and Grant, 2019, Føre et al., 2018, Antonucci and Costa, 2020), also known as smart aquaculture or precision fish farming, is a modern approach to fish and shellfish farming that integrates advanced technologies like IoT and data-driven techniques to optimize production efficiency, environmental sustainability, and animal welfare. It is often characterized by hundreds of interconnected sensors that store and serve data to gather essential information such as temperature, oxygen, weather, chlorophyll, fish tags, imagery data, and 3D sensor data, which can be utilized for monitoring fish variables (species, health, etc.), behavioral analysis, feeding decisions, and water quality prediction (Yang et al., 2021). Recently, DL has emerged as a promising solution to address the problems in precision aquaculture farming. For example, in Måløy et al. (2019), a dual-stream recurrent network was adopted to classify fish feeding status for salmon, yielding 80% feeding classification accuracy. However, these DL approaches rely heavily on large-scale and labeled datasets for training (Yang et al., 2021). Underwater conditions can be adverse, with poor illumination and low visibility in turbid water, as well as cluttered backgrounds, making it difficult to acquire high-quality data. Besides, it is very challenging to integrate, handle, and process numerous and varied data across multiple domains.
@@
Fortunately, FMs have become an encouraging solution to facilitate the realization of precision aquaculture farming, representing the unique capability to handle diverse data from multiple domains through few-short or zero-short learning. For example, generation-based FMs can be harnessed for data creation, while RLFMs can be utilized for decision-making in fish farming. The application of FMs could significantly enhance the accuracy and efficiency of associated applications in precision aquaculture farming.
3.3.5 Agricultural robots
@@
Agricultural robots, often referred to as “AgriRobots”, are at the forefront of transforming modern agriculture (Bechar and Vigneault, 2016). AgriRobots can automate various farming tasks such as planting, weeding, harvesting, and crop health monitoring (Fountas et al., 2020). For instance, AI-guided robots can identify and remove weeds with precision, significantly reducing the use of herbicides. In smart orchards, AI-guided robots equipped with a variety of sensors, including image sensors, LiDAR, and proximity sensors, autonomously pick ripe fruits, thereby optimizing harvesting efficiency and minimizing labor costs. Similarly, drones equipped with advanced imaging sensors can monitor fields for signs of disease or pests, enabling early intervention and reducing yield losses. Similar to the contexts of smart breeding and smart livestock farming, the development and implementation of AgriRobots require multidisciplinary knowledge including robotics, AI, plant science, environmental science, and data analytics. This mixed approach involves handling and processing a significant amount of data from various domains, making it a complex endeavor (Bechar and Vigneault, 2016, Fountas et al., 2020).
@@
The above challenges can be guided effectively by the potential application of FMs. These models, trained on large-scale multimodal data, can provide the robust predictive capabilities, which are necessary for advanced farming operations. For instance, RLFM (Section  2.2.4) could enable precision control of AgriRobots (Reed et al., 2022, Team et al., 2023a), while LFM (Section  2.2.1) could provide valuable insights during the design process of AgriRobots (Stella et al., 2023). By incorporating insights from diverse domains, FMs can offer an integrated solution, driving the development of more effective and efficient agricultural robots, and thus, revolutionizing the future of agriculture.
However, constructing an AFM from scratch presents significant challenges. Firstly, initial data collection requires a large and diverse dataset, which requires laborious efforts and potentially high costs (Lu and Young, 2020). The subsequent data labeling and curation require meticulousness and expertise. Additionally, the process of training models demands substantial computational resources, thus presenting impediments for individuals or entities with constrained resources. Lastly, fine-tuning the model for specific agricultural applications can be challenging, requiring a balance between generalizability and specificity. A detailed discussion of these challenges will be presented in Section  4.
@@ 4.1 Data collection
The process of data collection for training FMs in agricultural applications presents unique challenges due to the inherent variability and diversity in agriculture, influenced by factors such as crop types, growth stages, soil conditions, weather patterns, farming practices, etc. These factors result in a complex and highly diverse dataset, making its collection and standardization a challenging task (Lu and Young, 2020, Xu et al., 2023b). Firstly, it is critical to ensure the collected data is accurate, comprehensive, and adequately captures the wide variability found in different farming environments or diverse range of growth phases. However, acquiring high-quality data can be time-consuming, labor-intensive, and expensive, especially considering the need for ground truth labeling in supervised learning scenarios (Li et al., 2023a) and the variety of growth stages of plants or livestock. To tackle these challenges, generative models (Lu et al., 2022) that create high-fidelity images and label-efficient learning methods to save on tedious labeling costs (Li et al., 2023a, Zhang et al., 2023f) show promises. In addition, overcoming the time-consuming and labor-intensive data collection in the field can be achieved by leveraging advancements in autonomous data collection technologies, such as UAVs or UGVs equipped with sensors, as well as remote sensing techniques and the IoT devices (Mueller-Sim et al., 2017, del Cerro et al., 2021, Xu et al., 2022a). These technologies automate data-gathering processes and reduce the need for manual intervention. Secondly, data privacy and ownership pose challenges as farms are typically privately owned. Farmers may be reluctant to share data due to concerns over privacy or potential commercial exploitation. Novel frameworks, such as federated learning (Zhang et al., 2021a), can address these concerns, paving the way for broader access to high-quality, diverse agricultural datasets. Temporal dynamics represent another significant hurdle. Agricultural processes inherently change over time, influenced by daily fluctuations, seasonal variations, or annual changes. This requires the collection of time-series data (e.g., data across years), introducing another layer of complexity. To handle temporal dynamics, techniques such as time-series analysis (Ostrom, 1990), dynamic models based on Long Short-Term Memory (LSTM) (Graves and Graves, 2012), and data augmentation (Lu et al., 2022) can be employed to better account for the inherent temporal dynamics in agricultural data.